### 31 Jan 2024 v1

# Unimodular Gravity as an initial value problem

Abstract
Unimodular gravity is a compelling modified theory of gravity that has a natural solution to the cosmological constant problem. What is more, this theory has a generalized matter energy conservation law, when compared with that of general relativity. However, for unimodular gravity to be considered a viable alternative to general relativity, it has to have a well-posed initial value formulation. Working with vacuum unimodular gravity, we find the constraints required to have a consistent evolution and we show that, for such initial data, the evolution is well posed in the same way as general relativity. Then, we find sufficient conditions for a matter action to preserve the well-posedness of unimodular gravity. As a corollary, we reveal that the restriction on the spacetime volume, which gives rise to the particular properties of unimodular gravity, can be solved by an adequate choice of the lapse function.
I Introduction
General relativity (GR) is nowadays accepted as the theory of gravity. However, GR is not problem-free. In particular, it requires a cosmological constant to describe the universe at cosmological scales. The cosmological constant problem is the fact that the measured value for this constant departs, by many orders of magnitude, from that estimated by considering contributions of the fields in their vacuum states Weinberg (1989).
Unimodular gravity (UG) is a modified theory of gravity that could solve the cosmological constant problem by allowing for a cosmological constant to arise as an integration constant (for a critical review of this claim see Ref. Earman, 2022). Hence, UG has gained a lot of attention as a plausible theory of gravity that allows for matter energy-nonconservation Bonder et al. (2023).
One important property required of any fundamental physical theory is its ability to predict a system’s evolution from initial data. This evolution ought to be unique and it must leave the constraints valid. However, for some theories, these properties are not enough; the evolution must also be continuous and causal in the following sense: we expect that small perturbations on initial data produce small changes in the solutions, where the notion of smallness is given by Sobolev norms Wald (1984). In other words, we ask that the solutions depend continuously on the initial data, to avoid losing predictability by the fact that initial data can only be measured with finite precision. Also, changes in the initial data supported in a given region of the initial data hypersurface can only affect causally connected regions of spacetime. This idea is relevant in relativistic theories to have consistency with the spacetime causal structure. A theory in which the evolution is unique, continuous, causal, and such that it respects the constraints, is said to have a well-posed initial value formulation.
One of the most important features of GR is that it has a well-posed initial value formulation Wald (1984). This is not a trivial result since the metric, which is one of the evolving fields, has all the causal information, which is then used to check on the consistency of the evolution. Nevertheless, the well-posedness of the initial problem of GR can be proven by using the Bianchi identity to show that the constraints are kept valid under evolution and bringing the evolution equations, via a judicious choice of coordinates, to a form that allows one to prove that the above-mentioned properties hold. In addition, there exists a maximal development of the initial data Choquet-Bruhat and Geroch (1969).
It is worth mentioning that, besides GR, relatively few proofs of well-posed initial value formulation are known in modified gravity theories. Examples of modified gravity theories where such results have been obtained include scalar-tensor theories Salgado (2006), k-essence theory Rendall (2006), Einstein-æther theory Sarbach et al. (2019), Horndeski theories Kovács (2019); Kovács and Reall (2020), and a $4$-derivative scalar-tensor theory Saló et al. (2022). Still, to the best of our knowledge, there are no previous proofs of a proper initial value formulation for theories with nondynamical tensors.
The goal of this work is to show that UG has a well-posed initial value formulation. We must stress that the proof we present is not a simple application of the GR techniques since the UG constraints are different from those of GR. In this sense, this work presents relevant methods that may be used when investigating the initial value formulation of other modified gravity theories.
We structured the paper as follows: in Sec. II, to have a self-contained paper, we introduce the UG theory and the mathematical tools to study a relativistic theory of gravity in terms of $3$-dimensional geometrical objects that evolve. A reader familiar with these topics may skip this section. The most important part of this paper is the identification of the evolution and constraint equations of the theory, which is presented in Sec. III. In this section, we also perform the constraint analysis for UG and we carry out the analysis of the evolution equations, relying on the well-known BSSN formulation. Finally, we present our conclusions in Sec. IV.
II Preliminaries
II.1 Unimodular Gravity
Historically, UG dates back to Einstein Einstein (1919) and Pauli Pauli (1981) who were interested in possible interplays between gravity and elementary particles (for historical remarks see Ref. Álvarez and Velasco-Aja, 2023). Yet, the framework that is close to what is presented here emerged some fifty years later Anderson and Finkelstein (1971). Around that time the theory gained popularity when studied in the context of classical field theory van der Bij et al. (1982); Buchmuller and Dragon (1988, 1989). However, the current appeal of UG started with the observation that, in UG, the energy associated with the vacuum does not gravitate (in a semiclassical framework), bypassing the cosmological constant problem Ellis et al. (2011).
Recent works rekindled the interest in UG. For example, it has been shown that energy nonconservation avoids some incompatible features with Quantum Mechanics and could give rise to an effective cosmological constant that has an adequate sign and size Josset et al. (2017); Perez and Sudarsky (2019). Also, cosmological diffusion models in the UG framework affect the value of the Hubble constant Linares Cedeño and Nucamendi (2021), among other interesting features Linares Cedeño and Nucamendi (2023).
Here, we work on a $4$-dimensional spacetime $M$ equipped with the pseudo-Riemannian metric $g_{ab}$ (we follow the notation and conventions of Ref. Wald, 1984 and, in particular, pairs of indexes in between parenthesis/brackets stand for its symmetric/antisymmetric part with a $1/2$ factor). Moreover, spacetime is assumed to be globally hyperbolic, which allows us to foliate $M$ by constant time (Cauchy) hypersurfaces $\Sigma_{t}$.
There are several ways to introduce UG Carballo-Rubio et al. (2022); the UG action we consider here is
$\mathrm{S}\left[g^{ab},\lambda,\phi\right]$ $=$ $\frac{1}{2\kappa}\int\mathrm{d}^{4}x\left\{\sqrt{-g}R+\lambda\left(\sqrt{-g}-f\right)\right\}$ (1)
$+\mathrm{S_{M}}\left[g^{ab},\phi\right],$
where $g^{ab}$ is the inverse of $g_{ab}$, $\lambda$ is a scalar field that acts as a Lagrangian multiplier, $\kappa$ is the gravitational coupling constant, $R$ is the curvature scalar associated with the metric-compatible and torsion-free derivative $\nabla_{a}$ (for a study of UG with torsion see Ref. Bonder and Corral, 2018). Moreover, $g$ is the determinant of $g_{ab}$, and $f$ is a nondynamic positive scalar density, namely, a scalar that transforms under coordinate transformations mimicking $\sqrt{-g}$. Also, $\mathrm{S_{M}}$ is the matter action, which takes the form
$\mathrm{S_{M}}\left[g^{ab},\phi\right]\coloneqq\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{L}_{\mathrm{M}}\left(g^{ab},\phi\right),$ (2)
where $\mathcal{L}_{\mathrm{M}}\left(g^{ab},\phi\right)$ is the matter Lagrangian and $\phi$ collectively describes all the matter fields.
We emphasize that the only difference of the UG action when compared with that of GR is the presence of the term with the Lagrange multiplier. This term fixes the spacetime volume form $\sqrt{-g}\mathrm{d}^{4}x$ to coincide with $f\mathrm{d}^{4}x$, which explains the assumption that $f>0$. Notice that, by construction, the nondynamical function $f$ does not transform under (active) diffeomorphisms, partially “breaking” the invariance under diffeomorphisms. As a consequence, the UG matter conservation equation is more general than that of GR, as we argue below.
An arbitrary variation of the action (1) has the form
$\delta\mathrm{S}$ $=$ $\int\mathrm{d}^{4}x\left\{\sqrt{-g}\left[\frac{1}{2\kappa}(G_{ab}-\frac{1}{2}g_{ab}\lambda)-\frac{1}{2}T_{ab}\right]\delta g^{ab}\right.$ (3)
$+\left.\frac{1}{2\kappa}\left(\sqrt{-g}-f\right)\delta\lambda+\frac{\delta\mathcal{L}_{\mathrm{M}}}{\delta\phi}\delta\phi\right\},$
where we omit the boundary terms, something we do throughout the entire paper, and we define the energy-momentum tensor as
$T_{ab}\coloneqq-\frac{2}{\sqrt{-g}}\frac{\delta\left(\mathcal{L}_{\mathrm{M}}\sqrt{-g}\right)}{\delta g^{ab}}.$ (4)
Hence, the metric equation of motion is
$R_{ab}-\frac{1}{2}Rg_{ab}-\frac{1}{2}\lambda g_{ab}=\kappa T_{ab},$ (5)
Notice that $\lambda$ enters this last equation as a cosmological constant, even though at this point there is no reason for it to be constant. Furthermore, the equation of motion associated with $\lambda$ yields the “unimodular constraint”
$\sqrt{-g}=f.$ (6)
Of course, there are also matter field equations that can be written generically as
$\frac{\delta\mathcal{L}_{\mathrm{M}}}{\delta\phi}=0.$ (7)
The trace of Eq. (5) takes the form
$-2\lambda=\kappa T+R,$ (8)
where $T\coloneqq g^{ab}T_{ab}$ is the trace of the energy-momentum tensor. Introducing this trace into Eq. (5) yields
$R_{ab}-\frac{1}{4}Rg_{ab}=\kappa\left(T_{ab}-\frac{1}{4}Tg_{ab}\right),$ (9)
which is explicitly traceless. On the other hand, the divergence of Eq. (5) produces
$\kappa\nabla^{a}T_{ab}=-\frac{1}{2}\nabla_{b}\lambda,$ (10)
where we use the Bianchi identity. If $\nabla^{a}T_{ab}=0$, $\lambda$ becomes a constant, which acts as the cosmological constant. However, UG allows for more general matter solutions where $\nabla^{a}T_{ab}$ is not necessarily zero, opening the door to different phenomenology Bonder et al. (2023).
One of the main differences between GR and UG is the theories’ symmetries. As it is well known, GR is invariant under all diffeomorphisms, which in turn implies that $\nabla_{a}T^{ab}=0$. This is not the case in UG where the nondynamical function $f$ reduces invariance under diffeomorphisms, as it occurs whenever there are nondynamical fields Corral and Bonder (2019). To show this, consider the vacuum case $\mathrm{S_{M}}=0$. The variation of the vacuum UG action with respect to a diffeomorphism associated with $\xi^{a}$ is given by Eq. (3) with
$\delta g^{ab}$ $=$ £$\pounds_{\xi}g^{ab}=-2\nabla^{(a}\xi^{b)},$ (11)
$\delta\lambda$ $=$ £$\pounds_{\xi}\lambda=\xi^{a}\nabla_{a}\lambda.$ (12)
Then, the action variation takes the form
$\delta\mathrm{S}$ $=$ $\int\mathrm{d}^{4}x\left\{\left[\frac{1}{\kappa}(-G_{ab}+\frac{1}{2}g_{ab}\lambda)\right]\nabla^{a}\xi^{b}\right.$ (13)
$+\left.\frac{1}{2\kappa}\left(\sqrt{-g}-f\right)\xi^{a}\nabla_{a}\lambda\right\},$
where we use Eq. (7). After integrating by parts using the appropriate volume form and utilizing the Bianchi identity, this variation can be written as
$\delta\mathrm{S}$ $=$ $\frac{1}{2\kappa}\int\mathrm{d}^{4}x\sqrt{-g}\lambda\nabla_{a}\left(\xi^{a}\frac{f}{\sqrt{-g}}\right).$ (14)
On shell, Eq. (6) is valid, and thus, the vacuum action is only invariant under diffeomorphisms associated with divergence-free vector fields, namely, vector fields such that $\nabla_{a}\xi^{a}=0$. This restricted set of diffeomorphisms goes by the name of volume-preserving diffeomorphisms.
What is the matter conservation law associated with volume-preserving diffeomorphisms? To get an answer, we first notice that a divergence-free vector field $\xi^{a}$ can be written in terms of a generic tensor $\alpha_{ab}=-\alpha_{ba}$ as $\xi^{a}=\epsilon^{abcd}\nabla_{b}\alpha_{cd}$, where $\epsilon_{abcd}$ is the volume form associated with $g_{ab}$. Thus, the on-shell variation of $\mathrm{S_{M}}$ with respect to a volume-preserving diffeomorphism can be written as
$\delta\mathrm{S_{M}}$ $=$ $\int\mathrm{d}^{4}x\sqrt{-g}T_{ab}\nabla^{a}\epsilon^{bcde}\nabla_{c}\alpha_{de}.$ (15)
For the matter action to be invariant under any volume-preserving diffeomorphism, we require $\alpha_{de}$ to be arbitrary, which, after we integrate by parts, translates to
$\delta\mathrm{S_{M}}=\int d^{4}x\sqrt{-g}\alpha_{de}\epsilon^{bcde}\nabla_{c}\nabla^{a}T_{ab}.$ (16)
In turn, this implies that
$\nabla_{[b}\nabla^{a}T_{c]a}=0.$ (17)
This last equation is the UG matter conservation law, which is more general than that of GR. What is more, using the Poincaré Lemma Nakahara (2016) and under the hypothesis that $M$ is simply connected, which we assume, this conservation law implies that there exists a scalar $Q$ such that
$\nabla^{a}T_{ab}=\nabla_{b}Q.$ (18)
When $Q$ is constant, one recovers the GR matter conservation law. However, in UG, $Q$ can be arbitrary.
With this discussion, we cover all the aspects of UG necessary for our paper. We turn to present the framework to describe gravity theories in the framework of an evolving $3$-dimensional geometry.
II.2 Evolution in relativistic theories of gravity
Unlike other physical theories where spacetime is known and the main objective is to find the temporal evolution of other fields, in the context of relativistic gravitational theories, which includes GR and UG, spacetime is a priori unknown, and the goal is to deduce its geometry from initial data. In this section, we discuss fundamental aspects of the $3+1$-formalism we use to describe these theories by evolving geometrical objects.
Recall that we work under the assumption that $M$ can be foliated by Cauchy surfaces, $\Sigma_{t}$, parameterized by a global time function, $t$. We further assume that the normal vector $n^{a}$ to $\Sigma_{t}$ is timelike and is normalized according to $g_{ab}n^{a}n^{b}=-1$. On each hypersurface $\Sigma_{t}$, the spacetime metric induces a Riemann metric Wald (1984):
$h_{ab}=g_{ab}+n_{a}n_{b}.$ (19)
The inverse of this metric is $h^{ab}=g^{ab}+n^{a}n^{b}$, where $g_{ab}$ is the only metric used throughout the text to raise/lower indices. Additionally, $h_{ab}n^{a}=0$, which can be corroborated by using Eq. (19). What is more, $\tensor{h}{{}^{a}_{b}}$ is a projector onto $\Sigma_{t}$.
Let $t^{a}$ be a timelike vector field in $M$ defined by
$t^{a}\nabla_{a}t=1.$ (20)
This vector field identifies points on infinitesimally close hypersurfaces of constant $t$, giving a flow of time. We can decompose $t^{a}$ into its normal and tangential parts as
$t^{a}=Nn^{a}+N^{a},$ (21)
where $N$ is the lapse function and $N^{a}$ is the shift vector. Relevantly, the fact that $t^{a}$ must be timelike implies that $N>0$. Broadly speaking, $N$ gives the rate of change of physical time as compared with $t$. On the other hand, $N^{a}$ tells us how the coordinates are transported from a point in $\Sigma_{t}$ to a point in $\Sigma_{t+\mathrm{d}t}$. Fig. 1 illustrates this construction.
Figure 1: A spacetime diagram illustrating the definition of the field $t^{a}$, the shift vector, $N^{a}$, and the lapse function, $N$.
We can find the covariant derivative associated with $h_{ab}$. First, consider a spacetime vector field $v^{a}$. We can express this field as
$v^{a}=v_{\perp}n^{a}+v^{a}_{\parallel},$ (22)
where $v^{a}_{\parallel}$ is the component parallel to $\Sigma_{t}$. Note that, when $v^{a}=v^{a}_{\parallel}$, we can think of $v^{a}$, restricted on $\Sigma_{t}$, as a vector field on $\Sigma_{t}$. More generally, a spacetime tensor $\tau^{a_{1}\cdots a_{k}}_{b_{1}\cdots b_{l}}$ is tangent to $\Sigma_{t}$ if
$\tau^{a_{1}\cdots a_{k}}_{b_{1}\cdots b_{l}}=\tensor{h}{{}^{a_{1}}_{c_{1}}}\cdots\tensor{h}{{}^{a_{k}}_{c_{k}}}\tensor{h}{{}_{b_{1}}^{d_{1}}}\cdots\tensor{h}{{}_{b_{l}}^{d_{l}}}\tau^{c_{1}\cdots c_{k}}_{d_{1}\cdots d_{l}}.$ (23)
We define the derivative operator on $\Sigma_{t}$, $D_{a}$, by
$D_{e}\tau^{a_{1}\cdots a_{k}}_{b_{1}\cdots b_{l}}=\tensor{h}{{}^{a_{1}}_{c_{1}}}\cdots\tensor{h}{{}^{a_{k}}_{c_{k}}}\tensor{h}{{}_{b_{1}}^{d_{1}}}\cdots\tensor{h}{{}_{b_{l}}^{d_{l}}}\tensor{h}{{}_{e}^{f}}\nabla_{f}\tau^{c_{1}\cdots c_{k}}_{d_{1}\cdots d_{l}},$ (24)
where $\tau^{a_{1}\cdots a_{k}}_{b_{1}\cdots b_{l}}$ is a tangential tensor. Importantly, one can readily show that this is the only (torsionless) derivative operator such that $D_{a}h_{bc}=0$ Wald (1984).
We turn now to introduce the time derivative of any tangential tensor, denoted with a dot. This derivative is defined by
$\dot{\tau}^{a_{1}\cdots a_{k}}_{b_{1}\cdots b_{l}}=\tensor{h}{{}^{a_{1}}_{c_{1}}}\cdots\tensor{h}{{}^{a_{k}}_{c_{k}}}\tensor{h}{{}_{b_{1}}^{d_{1}}}\cdots\tensor{h}{{}_{b_{l}}^{d_{l}}}\pounds_{t}\tau^{c_{1}\cdots c_{k}}_{d_{1}\cdots d_{l}},$ (25)
which is, by construction, a tangential tensor, and where £$\pounds_{t}$ stands for the Lie derivative along $t^{a}$.
One of the dynamical variables we use to describe the evolution of spacetime’s geometry is $h_{ab}$; thus, we need to calculate its time derivative. We define
$K_{ab}\coloneqq\tensor{h}{{}_{a}^{c}}\nabla_{c}n_{b},$ (26)
which can be shown to be symmetric. The tensor $K_{ab}$ is known as the extrinsic curvature, and it describes the embedding of $\Sigma_{t}$ in $M$. If we use the time derivative on $h_{ab}$, we obtain
$\dot{h}_{ab}=2NK_{ab}+2D_{(a}N_{b)},$ (27)
showing that $K_{ab}$ is related to the time derivative of $h_{ab}$. Hence, and given that the equations of motion for UG are second order, it is natural to propose that the appropriate initial data for UG should be given by $h_{ab}$ and $K_{ab}$.
To end this subsection, we state certain expressions that relate the Riemann tensor associated with $g_{ab}$, $\tensor{R}{{}_{a}{}_{b}{}_{c}^{d}}$, with objects in $\Sigma_{t}$. One can show that the purely tangential projection of the Riemann tensor satisfies a Gauss-Codazzi relation Wald (1984) of the form
$\tensor{R}{{}_{a}{}_{b}{}_{c}^{d}}\tensor{h}{{}^{a}_{e}}\tensor{h}{{}^{b}_{f}}\tensor{h}{{}^{c}_{g}}\tensor{h}{{}_{d}^{j}}=\tensor{{}^{\left(3\right)}R}{{}_{e}{}_{f}{}_{g}^{j}}+K_{eg}\tensor{K}{{}_{f}^{j}}-K_{fg}\tensor{K}{{}_{e}^{j}},$ (28)
where $\tensor{{}^{\left(3\right)}R}{{}_{a}{}_{b}{}_{c}^{d}}$ is the Riemann tensor associated with $h_{ab}$. In addition, $\tensor{{}^{\left(3\right)}R}{{}_{a}{}_{b}}$ and ${}^{\left(3\right)}R$ respectively represent the $3$-dimensional Ricci tensor and curvature scalar. Another projection of the Riemann tensor satisfies
$\tensor{R}{{}_{a}{}_{b}{}_{c}^{d}}\tensor{h}{{}^{a}_{e}}\tensor{h}{{}^{b}_{f}}\tensor{h}{{}^{c}_{g}}n_{d}=D_{e}K_{fg}-D_{f}K_{eg}.$ (29)
We can also prove that
$\tensor{R}{{}_{a}{}_{b}{}_{c}^{d}}n^{a}\tensor{h}{{}^{b}_{e}}\tensor{h}{{}^{c}_{f}}n_{d}$ $=$ $\tensor{h}{{}_{e}^{b}}\tensor{h}{{}_{f}^{c}}n^{a}\nabla_{a}K_{bc}-a_{e}a_{f}$ (30)
$-D_{\left(e\right.}a_{\left.f\right)}+\tensor{K}{{}_{e}^{a}}K_{af},$
where $a^{a}\coloneqq n^{b}\nabla_{b}n^{a}$ is a tangential vector field.
Other useful projections are:
$R_{cd}\tensor{h}{{}^{c}_{a}}\tensor{h}{{}^{d}_{b}}$ $=$ ${{}^{\left(3\right)}R_{ab}}+KK_{ab}-2\tensor{K}{{}_{a}^{c}}K_{cb}+N^{-1}\dot{K}_{ab}$ (31)
$-N^{-1}N^{c}D_{c}K_{ab}-2N^{-1}K_{c\left(a\right.}D_{\left.b\right)}N^{c}$
$-a_{a}a_{b}-D_{\left(a\right.}a_{\left.b\right)},$
$R_{cd}n^{c}n^{d}$ $=$ $-N^{-1}h^{cd}\dot{K}_{cd}+N^{-1}N^{c}D_{c}K+a_{c}a^{c}$
$+2N^{-1}K^{cd}D_{c}N_{d}+K^{cd}K_{cd}+D_{c}a^{c},$
$R_{bc}n^{b}\tensor{h}{{}^{c}_{a}}$ $=$ $D_{b}\tensor{K}{{}^{b}_{a}}-D_{a}K,$ (33)
$R$ $=$ ${{}^{\left(3\right)}}R+K^{2}-3K_{cd}K^{cd}+2N^{-1}h^{cd}\dot{K}_{cd}$ (34)
$-2N^{-1}N^{c}D_{c}K-4N^{-1}K^{cd}D_{c}N_{d}$
$-2a_{c}a^{c}-2D_{c}a^{c},$
where $K\coloneqq\tensor{K}{{}_{a}^{a}}$.
Finally, we can show that
$\dot{K}_{ef}$ $=$ $N\tensor{h}{{}_{e}^{b}}\tensor{h}{{}_{f}^{c}}n^{a}\nabla_{a}K_{bc}+N^{a}D_{a}K_{ef}$ (35)
$+2K_{a\left(e\right.}D_{\left.f\right)}N^{a}+2N\tensor{K}{{}_{e}^{a}}K_{af}.$
With this result, Eq. (30) can be written in a form that is useful below:
$\tensor{R}{{}_{a}{}_{b}{}_{c}^{d}}n^{a}\tensor{h}{{}^{b}_{e}}\tensor{h}{{}^{c}_{f}}n_{d}$ $=$ $N^{-1}\dot{K}_{ef}-N^{-1}N^{a}D_{a}K_{ef}$ (36)
$-2N^{-1}K_{a\left(e\right.}D_{\left.f\right)}N^{a}-\tensor{K}{{}_{e}^{a}}K_{af}$
$-a_{e}a_{f}-D_{\left(e\right.}a_{\left.f\right)}.$
These are all the technical tools we require. In the next section, we classify the UG equations of motion as evolution or constraints, and we find the additional conditions for the constraints to be kept valid under evolution.
III Initial value problem
The first task when studying the initial value problem of a theory is to identify the constraints. In this section, we classify the UG field equations into evolution equations and constraints. When working with a geometrical gravity theory with second-order equations of motion, an evolution equation, by definition, contains time derivatives of the extrinsic curvature, which can be thought of as second-time derivatives of $h_{ab}$. Conversely, an equation with no time derivatives of $K_{ab}$ is a constraint, which must be imposed on the initial data, and, for consistency, must be kept valid under evolution. Notice that we cannot perform the initial value study for generic matter fields without specifying the matter action. Thus, in what follows, only vacuum UG is considered; a set of sufficient conditions for the matter equations of motion to not affect the well-posedness of the UG initial value problem is given in Subsec. III.3.
We define
$E_{ab}\coloneqq R_{ab}-\frac{1}{4}Rg_{ab}.$ (37)
The vanishing of this tensor represents the vacuum UG equations of motion, Eq. (9), which, in this study, are not assumed to hold throughout $M$. Instead, a weaker assumption is considered: the constraints are only valid on the initial data hypersurface, while the evolution equation is satisfied in $M$. Interestingly, we will prove that dynamical consistency implies that $E_{ab}=0$ in $M$.
Moreover, using the Bianchi identity, we can show that
$\nabla^{a}E_{ab}=\frac{1}{4}\nabla_{b}R.$ (38)
We emphasize that, since we do not assume that $E_{ab}$ vanishes on $M$, we cannot claim that its divergence also vanishes. The vanishing of $\nabla^{a}E_{ab}$ in $M$ immediately implies, from Eq. (38), that $R$ is constant, something we cannot show at this stage.
Notice that the unimodular constraint, Eq. (6), is a constraint in the sense that it does not have time derivatives of $K_{ab}$. Still, it is not a relation that can be imposed on the initial data and be automatically valid throughout spacetime by the evolution since the values of $f$, which is nondynamical, are given a priori all over $M$. Thus, we do not verify this property on the unimodular constraint. One of the main results of this paper is that the unimodular constraint can be satisfied by choosing the lapse function.
Also, $E_{ab}$ is such that $g^{ab}E_{ab}=0$. Thus, we can show that
$h^{ab}E_{ab}=E_{ab}n^{a}n^{b}.$ (39)
Eq. (39) allows us to identify the $3$-dimensional trace of tangential-tangential projection with the normal-normal projection, rendering the latter redundant. Therefore, in what follows, we omit the normal-normal projection. In addition, there are only nine independent components of $E_{ab}=0$. These components, together with the unimodular constraint, amount to ten equations, which coincides with the number of independent equations of GR.
We separate $E_{ab}=0$ into its tangential-tangential projection and its normal-tangential projection. The tangential-tangential projection and normal-tangential projection of $E_{ab}=0$ are respectively given by
00 0 $=$ ${{}^{\left(3\right)}R_{ab}}+KK_{ab}-2\tensor{K}{{}_{a}^{c}}K_{cb}+N^{-1}\dot{K}_{ab}-a_{a}a_{b}$ (40)
$-N^{-1}N^{c}D_{c}K_{ab}-2N^{-1}K_{c(a}D_{b)}N^{c}-D_{(a}a_{b)}$
$-\frac{1}{4}\big{[}{{}^{\left(3\right)}}R+K^{2}-3K_{cd}K^{cd}+2N^{-1}h^{cd}\dot{K}_{cd}$
$-2N^{-1}N^{c}D_{c}K-4N^{-1}K^{cd}D_{c}N_{d}-2a_{c}a^{c}$
$-2D_{c}a^{c}\big{]}h_{ab},$
and
$0=D_{b}\tensor{K}{{}^{b}_{a}}-D_{a}K.$ (41)
where we use Eqs. (31) and (33), and we define the tangential tensor $\mathcal{E}_{ab}\coloneqq E_{cd}\tensor{h}{{}^{c}_{a}}\tensor{h}{{}^{d}_{b}}$. Equation (40) has time derivatives of $K_{ab}$. Therefore, the tangential-tangential projection of the field equation is an evolution equation. On the other hand, the normal-tangential projection, Eq. (41) does not contain time derivatives of $K_{ab}$, and it is thus a constraint equation that is reminiscent of the momentum constraint in GR.
In summary, in terms of components, six of the field equations of vacuum UG, the tangent-tangent projections, are evolution equations, while the remaining three, the normal-tangential projections, are constraints. In the next subsection, we analyze the evolution of the constraints.
III.1 Constraint equations
We now study if the constraint given in Eq. (41) is maintained under evolution. In other words, given initial data on $\Sigma_{t}$ that satisfies Eq. (41), we must check if the fields obtained by evolving this initial data still satisfy the constraint. This can be shown by proving that the constraints on the initial data hypersurface have zero time derivative. Without loss of generality, we take the initial value hypersurface to coincide with the $t=0$ surface, $\Sigma_{0}$.
We define
$C_{a}\coloneqq D_{b}\tensor{K}{{}^{b}_{a}}-D_{a}K.$ (42)
This tangential tensor is such that, when it vanishes, the constraints are satisfied. What is more, since we are considering the evolution of the dynamical fields, the evolution equations, Eq. (40), are assumed to be valid throughout $M$. Let $A_{1}$ be the following set of assumptions: $C_{a}=0$ on $\Sigma_{0}$ and $\mathcal{E}_{ab}=0$ in $M$. What remains to check is if, under $A_{1}$, $\dot{C}_{a}=0$ on $\Sigma_{0}$.
From the definition of the time derivative, we readily obtain
$\dot{C}_{a}=N\tensor{h}{{}_{a}^{d}}n^{e}\nabla_{e}C_{d}+N^{e}D_{e}C_{a}+NC_{e}\tensor{K}{{}^{e}_{a}}+C_{e}D_{a}N^{e}.$ (43)
Given that $C_{a}=0=D_{a}C_{b}$ on $\Sigma_{0}$, Eq. (43), when restricted to $\Sigma_{0}$, takes the form
$\left.\dot{C}_{a}\right|_{\Sigma_{0}}=N\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}C_{b},$ (44)
which is not automatically zero under $A_{1}$.
Let
$S_{a}\coloneqq\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}C_{b}.$ (45)
Since $N>0$, the time derivative of $C_{a}$ vanishes on $\Sigma_{0}$ if $S_{a}=0$ on $\Sigma_{0}$. We can show that, under $A_{1}$, $S_{a}=-D_{a}R/4$ on $\Sigma_{0}$:
Proof.
We can readily verify that
$\tensor{h}{{}_{a}^{d}}\nabla^{b}\left(E_{bc}\tensor{h}{{}^{c}_{d}}\right)=D^{b}\mathcal{E}_{ab}+\mathcal{E}_{ab}a^{b}-S_{a}-KC_{a},$ (46)
where we use $K=\nabla_{c}n^{c}$.
Now, using Eq. (38), we obtain
$\tensor{h}{{}_{a}^{d}}\nabla^{b}\left(E_{bc}\tensor{h}{{}^{c}_{d}}\right)=\frac{1}{4}D_{a}R+C_{b}\tensor{K}{{}_{a}^{b}}-a_{a}\mathcal{E},$ (47)
where
$\mathcal{E}$$\coloneqq$ $\mathcal{E}_{ab}h^{ab}$ (48)
$=$ $\frac{1}{4}{{}^{\left(3\right)}R}+\frac{1}{4}K^{2}+\frac{1}{4}K_{ab}K^{ab}-\frac{1}{2}N^{-1}h^{ab}\dot{K}_{ab}$
$+\frac{1}{2}a_{a}a^{a}+\frac{1}{2}N^{-1}N^{c}D_{c}K+N^{-1}K^{ab}D_{a}N_{b}$
$+\frac{1}{2}D_{a}a^{a}.$
Combining Eqs. (46) and (47) produces
$S_{a}=-\frac{1}{4}D_{a}R-\tensor{K}{{}_{a}^{b}}C_{b}-KC_{a}+\mathcal{E}_{ab}a^{b}+a_{a}\mathcal{E}+D^{b}\mathcal{E}_{ab}.$ (49)
Hence, on $\Sigma_{0}$ and assuming $A_{1}$,
$\left.S_{a}\right|_{\Sigma_{0}}=-\frac{1}{4}\left.D_{a}R\right|_{\Sigma_{0}}.$ (50)
According to Dirac’s method to study theories with constraints Dirac (2001), it is necessary to promote $S_{a}=0$ as a (secondary) constraint. This, of course, implies that the $4$-dimensional curvature scalar, $R$, must be constant throughout $\Sigma_{0}$. Namely, $R=4\Lambda_{0}$, where $D_{a}\Lambda_{0}=0$. We consider it necessary to make two comments: first, we must consider $R$ as a shorthand notation for the right-hand side of Eq. (34), which only contains tangential objects. Second, as the notation suggests, $\Lambda_{0}$ will end up playing the role of the cosmological constant. However, at this stage, we can only claim that is constant along $\Sigma_{0}$ and there is no reason to assume that $\dot{\Lambda}_{0}=0$. dynamical consistency will require $\Lambda_{0}$ to be constant throughout spacetime, as we show next.
We now prove that, assuming $A_{1}$, $\dot{R}=0$ on $\Sigma_{0}$:
Proof.
Using the fact that $\tensor{h}{{}_{a}^{b}}-n_{a}n^{b}$ is the identity tensor in spacetime, we can verify that
$E_{ab}n^{b}=C_{a}-n_{a}\mathcal{E}.$ (51)
The divergence of this last equation produces
$\nabla^{a}(E_{ab}n^{b})=\nabla_{a}C^{a}-K\mathcal{E}-n^{a}\nabla_{a}\mathcal{E}.$ (52)
We can calculate the same object using the Leibniz rule and Eq. (38). This calculation yields
$\nabla^{a}\left(E_{ab}n^{b}\right)=\frac{1}{4}n^{a}\nabla_{a}R+\mathcal{E}_{ab}K^{ab}-C_{a}a^{a}.$ (53)
When we compare these results, we obtain
$D_{a}C^{a}+2C_{a}a^{a}-K\mathcal{E}-n^{a}\nabla_{a}\mathcal{E}=\frac{1}{4}n^{a}\nabla_{a}R+\mathcal{E}_{ab}K^{ab}.$ (54)
where we use $\nabla_{a}C^{a}=D_{a}C^{a}+C_{a}a^{a}$, which follows directly from the definition of $D_{a}$. Hence, on $\Sigma_{0}$ and assuming $A_{1}$,
$\left.n^{a}\nabla_{a}R\right|_{\Sigma_{0}}=0.$ (55)
This result, together with the fact that $D_{a}R=0$ on $\Sigma_{0}$, allows us to prove that $\dot{R}=0$ on $\Sigma_{0}$. ∎
The lesson from the last proof is that, under $A_{1}$, $\Lambda_{0}$ is constant in $M$. Recall that $A_{1}$ is a significantly weaker set of assumptions than assuming that $E_{ab}=0$ in $M$.
We now need to check if $\dot{S}_{a}=0$ on $\Sigma_{0}$, assuming $A_{1}$ and $\left.S_{a}\right|_{\Sigma_{0}}=0$; we refer to this new set of assumptions by $A_{2}$. As expected, under $A_{2}$, the vanishing of $\dot{S}_{a}$ is equivalent to
$\left.\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}S_{b}\right|_{\Sigma_{0}}=0,$ (56)
which we show to hold:
Proof.
Using Eq. (49), we get
$\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}S_{b}$ $=$ $-\frac{1}{4}\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}D_{b}R-C_{d}\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}\tensor{K}{{}_{b}^{d}}$ (57)
$-\tensor{K}{{}_{a}^{b}}S_{b}-C_{a}n^{b}\nabla_{b}K-KS_{a}$
$+a^{d}\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}\mathcal{E}_{bd}+\mathcal{E}_{ab}n^{c}\nabla_{c}a^{b}+a_{a}n^{b}\nabla_{b}\mathcal{E}$
$+\mathcal{E}\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}a_{b}+\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}D^{d}\mathcal{E}_{bd},$
where we write $n^{c}\nabla_{c}\tensor{K}{{}_{a}{}_{b}}$ and $n^{a}\nabla_{a}K$ in terms of $R=4\Lambda_{0}$ and other tangential objects using Eqs. (34) and (35). Notice that all terms but the first are proportional to $C_{a}$, $S_{a}$, $\mathcal{E}$, $\mathcal{E}_{ab}$ or derivatives of $\mathcal{E}$ and $\mathcal{E}_{ab}$, all of which vanish under $A_{2}$.
We only need to show that $\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}D_{b}R$ vanishes under $A_{2}$. For this, we use
$\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}D_{b}R$ $=$ $\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}(\tensor{h}{{}_{b}^{d}}\nabla_{d}R)$
$=$ $\tensor{h}{{}_{a}^{b}}(n^{c}\nabla_{c}\tensor{h}{{}_{b}^{d}})\nabla_{d}R+\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}\nabla_{b}R$
$=$ $a_{a}n^{b}\nabla_{b}R+\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{b}\nabla_{c}R$
$=$ $a_{a}n^{b}\nabla_{b}R+D_{a}(n^{b}\nabla_{b}R)-\tensor{K}{{}_{a}^{b}}D_{b}R.$
Here, we repeatedly use the definition of the tangential derivative, the Leibniz rule, and the fact that covariant derivatives acting on scalars commute; this is a consequence of the torsion-free hypothesis (a study of an unimodular theory with torsion is presented in Ref. Bonder and Corral, 2018). Upon evaluation on $\Sigma_{0}$, the first and last terms in Eq. (LABEL:42) can be seen to vanish using $R=4\Lambda_{0}$. Moreover, if we take the tangential derivative of Eq. (54), we can write the second term in terms of objects that also vanish under $A_{2}$. With all this, it is clear that under $A_{2}$, $\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}S_{b}=0$ on $\Sigma_{0}$. ∎
The main conclusion of this subsection is that, for consistency, the initial data in vacuum UG must be such that $C_{a}=0$ and $S_{a}=0$. Note that we can enforce $S_{a}=0$ by choosing initial data such that $R$ is constant on $\Sigma_{0}$. As a consequence, we get $R=4\Lambda_{0}$ out of a weaker set of assumptions. In the next subsection, we verify that the evolution of the system is unique, continuous, and causal in the sense described above.
III.2 Evolution equations
In the previous section, we show that the constraints are maintained under the temporal evolution, provided that the initial data satisfies $C_{a}=0$ and $S_{a}=0$. In this section, we study the evolution equations, namely, the tangential-tangential projection of the field equations, and the unimodular constraint.
We have taken $h_{ab}$ and $K_{ab}$ as the dynamic variables for vacuum UG. However, an approach that better adapts to our needs is the BSSN formulation, developed for GR initially by Shibata and Nakamura Shibata and Nakamura (1995) and later by Baumgarte and Shapiro Baumgarte and Shapiro (1998). The BSSN formulation consists of separating the conformal factor for the spatial metric and the trace of the extrinsic curvature and then studying their evolution separately. In this subsection, we consider components on the Cauchy surfaces of all the tensors, which are denoted with Latin indexes $i,j,k,\ldots$. Also, for simplicity, we take the shift vector as $N^{a}=0$, which can be trivially relaxed.
Let us consider the conformal transformation of the spatial metric:
$h_{ij}=\psi^{4}\tilde{h}_{ij},$ (59)
where $\psi$ is fixed by $\tilde{h}=1$, where $\tilde{h}$ is the determinant of $\tilde{h}_{ij}$. Hence,
$\psi=h^{1/12},$ (60)
with $h$ the determinant associated with $h_{ij}$. Thus, we can rewrite Eq. (59) as
$\tilde{h}_{ij}=h^{-1/3}h_{ij}.$ (61)
From Eqs. (59) and (61), it is easy to see that the inverse spatial metric is given by
$\tilde{h}^{ij}=\psi^{4}h^{ij}=h^{1/3}h^{ij}.$ (62)
On the other hand, let $A_{ij}$ be the symmetric tensor field defined by
$A_{ij}\coloneqq K_{ij}-\frac{1}{3}Kh_{ij}.$ (63)
Note that this tensor is the traceless part of the extrinsic curvature, and thus, $A_{ij}\tilde{h}^{ij}=0$.
In the BSSN formulation, instead of using $h_{ab}$ and $K_{ab}$ as the dynamical variables, we use $\phi$, $K$, $\tilde{h}_{ab}$, $\tilde{A}_{ab}$ and $\tilde{\Gamma}^{a}$, which are given by
$\phi$$\coloneqq$ $\ln\psi=\frac{1}{12}\ln h,$ (64a)
$K$ $=$ $h^{ij}K_{ij},$ (64b)
$\tilde{h}_{ij}$ $=$ $\mathrm{e}^{-4\phi}h_{ij},$ (64c)
$\tilde{A}_{ij}$ $=$ $\mathrm{e}^{-4\phi}A_{ij},$ (64d)
$\tilde{\Gamma}^{i}$$\coloneqq$ $\tilde{h}^{jk}\ \tensor{\tilde{\Gamma}}{{}^{i}_{j}{}_{k}}.$ (64e)
The first expression is a redefinition of the conformal factor, the second is the trace of the extrinsic curvature, the third is the conformal transformation given in Eq. (59), the fourth is a rescaling of the traceless part of the extrinsic curvature, and finally, $\tilde{\Gamma}^{i}$ are the conformal connection functions, where
$\tensor{\tilde{\Gamma}}{{}^{i}_{j}{}_{k}}=\frac{1}{2}\tilde{h}^{il}\left(-\partial_{l}\tilde{h}_{jk}+\partial_{j}\tilde{h}_{kl}+\partial_{k}\tilde{h}_{lj}\right),$ (65)
are the Christoffel symbols associated with $\tilde{h}_{ij}$. Furthermore, from fact that $\tilde{h}=1$, we can easily show that Eq. (64e) is equivalent to
$\tilde{\Gamma}^{a}=-\partial_{j}\tilde{h}^{ji}.$ (66)
What remains to be done is to rewrite the constraint and evolution equations in terms of the BSSN variables; this allows us to argue when comparing with the GR evolution equations, that vacuum UG has a well-posed initial value formulation.
Given that $R=4\Lambda_{0}$ and $\mathcal{E}=0$, where $\Lambda_{0}$ is constant, and using Eqs. (34) and (48), we can show that
$4\Lambda_{0}=2{{}^{\left(3\right)}}R+2K^{2}-2K_{cd}K^{cd},$ (67)
where the right-hand side is reminiscent of the GR Hamiltonian constraint. Thus, we can cast the tangential-tangential projection of the UG field equation, Eq. (40), as
00 0 $=$ ${{}^{\left(3\right)}}R_{ab}+KK_{ab}-2\tensor{K}{{}_{a}^{c}}K_{cb}+N^{-1}\dot{K}_{ab}$ (68)
$-N^{-1}N^{c}D_{c}K_{ab}-2N^{-1}K_{c\left(a\right.}D_{\left.b\right)}N^{c}$
$-N^{-1}D_{a}D_{b}N-\Lambda_{0}h_{ab},$
which coincides with the evolution equation for $\dot{K}_{ab}$ obtained in GR with a cosmological constant $\Lambda_{0}$. Moreover, all other evolution equations are identical to those of GR. Thus, in vacuum UG, the system of evolution equations, in BSSN variables, takes the form
$\partial_{t}\tilde{h}_{ij}$ $=$ $2N\tilde{A}_{ij},$ (69a)
$\partial_{t}\phi$ $=$ $\frac{1}{6}NK,$ (69b)
$\partial_{t}\tilde{A}_{ij}$ $=$ $\mathrm{e}^{-4\phi}\left(D_{i}D_{j}N-N{{}^{\left(3\right)}}R_{ij}\right)^{\mathrm{TL}}-$ (69c)
$N\left(K\tilde{A}_{ij}-2\tensor{\tilde{A}}{{}_{i}^{k}}\tilde{A}_{kj}\right),$
$\partial_{t}K$ $=$ $D_{k}D^{k}N-N\left(\tilde{A}_{ij}\tilde{A}^{ij}+\frac{1}{3}K^{2}-\Lambda_{0}\right),$ (69d)
$\partial_{t}\tilde{\Gamma}^{i}$ $=$ $-2N\left(\tensor{\tilde{\Gamma}}{{}^{i}_{j}{}_{k}}\tilde{A}^{jk}+6\tilde{A}^{ij}\partial_{j}\phi-\frac{2}{3}\tilde{h}^{ij}\partial_{j}K\right)$ (69e)
$+2\tilde{A}^{ij}\partial_{j}N,$
which is completely equivalent to that of GR with a cosmological constant. Given that GR with a cosmological constant has a well-posed initial value formulation (see, e.g., Sarbach et al. (2002) and recall that the cosmological constant term does not have metric derivatives), we can conclude that vacuum UG also has a well-posed initial value formulation.
We turn now to discuss the unimodular constraint. Using the well-known expression $\sqrt{-g}=N\sqrt{h}$ and Eq. (64a), we get $\sqrt{-g}=N\mathrm{e}^{6\phi}$. On the other hand, the unimodular constraint is $\sqrt{-g}=f$. By direct comparison, we can conclude that the unimodular constraint is satisfied as long as
$N=f\mathrm{e}^{-6\phi}.$ (70)
Notably, $f>0$, avoiding the possibility that we have an ill-defined lapse function. Moreover, the fact that the unimodular constraint can be solved by simply choosing the lapse function is compatible with the result of Ref. Bengochea et al. (2023) where it is noted that one component of the metric is sufficient to solve the unimodular constraint.
With this analysis, we shown that, in vacuum UG, the evolution of the initial data is unique, continuous, and causal in a neighborhood associated with the initial data. However, there are cases in which the initial data can only be given on different charts covering $\Sigma_{0}$. In this case, a “gluing” of the different evolutions must be made. Fortunately, this procedure can be carried out in the same manner as in GR Wald (1984), since this procedure relies on making coordinate transformations that are available in UG. Likewise, there is no obstruction to applying the GR proof Choquet-Bruhat and Geroch (1969) that shows that there is a maximal evolution of the initial data. Therefore, we can conclude that vacuum UG has a well-posed initial value formulation which gives rise to a maximal evolution of the initial data. In the next subsection, we discuss the introduction of matter.
III.3 Initial value formulation with matter
The initial value formulation of a given UG theory with matter can only be studied properly once the matter action is given. Still, we provide a set of sufficient conditions to have a well-posed initial value formulation for a UG theory with a generic matter action. The matter UG field equation associated with the metric, Eq. (9), is obtained when
$E_{ab}^{\rm Matter}\coloneqq R_{ab}-\frac{1}{4}Rg_{ab}-\kappa\left(T_{ab}-\frac{1}{4}Tg_{ab}\right),$ (71)
vanishes. Again, we do not assume that his object vanishes $M$ but only a weaker assumption: the constraints vanish on $\Sigma_{0}$ and the evolution equations do vanish in $M$. Importantly, we can check that
$\nabla^{a}E_{ab}^{\rm Matter}=\frac{1}{4}\nabla_{b}\left(R-4\kappa Q+\kappa T\right).$ (72)
where we use Eq. (18).
The tangential-tangential projection and normal-tangential projection of Eq. (71) are respectively given by
00 0 $=$ $R_{cd}\tensor{h}{{}^{c}_{a}}\tensor{h}{{}^{d}_{b}}-\frac{1}{4}Rh_{ab}-\kappa\left(T_{cd}\tensor{h}{{}^{c}_{a}}\tensor{h}{{}^{d}_{b}}-\frac{1}{4}Th_{cd}\right),$
00 0 $=$ $R_{bc}n^{b}\tensor{h}{{}^{c}_{a}}-\kappa T_{bc}n^{b}\tensor{h}{{}^{c}_{a}},$ (74)
where $\mathcal{E}^{\rm Matter}_{ab}\coloneqq E^{\rm Matter}_{cd}\tensor{h}{{}^{c}_{a}}\tensor{h}{{}^{d}_{b}}$, which is a tangential tensor. Note that, since $E_{ab}^{\rm Matter}g_{ab}=0$, the normal-normal projection of Eq. (71) can be obtained from Eq. (III.3), and thus, we do not consider it, as before.
Repeating the analysis of vacuum UG, we can conclude that Eq. (III.3) is an evolution equation. On the other hand, the normal-tangential projection, Eq. (74), takes the form
$0=D_{b}\tensor{K}{{}^{b}_{a}}-D_{a}K-\kappa T_{bc}n^{b}\tensor{h}{{}^{c}_{a}},$ (75)
which is a constraint provided that $T_{bc}n^{b}\tensor{h}{{}^{c}_{a}}$ does not contain second time derivatives of $h_{ab}$ or the matter fields. Assuming this is the case, we verify if this constraint is maintained under temporal evolution.
The analysis of the consistency of the constraints can be done in the same way as that presented in the Subsec. III.1 for vacuum UG. The only difference is that what is constant in the case where there is matter is $\Lambda_{0}=R-4\kappa Q+\kappa T$, as expected from Eq. (72). The conclusion is then that, to have dynamical consistency, the initial data must be subject to
$C_{a}^{\rm Matter}$$\coloneqq$ $D_{b}\tensor{K}{{}^{b}_{a}}-D_{a}K-\kappa T_{bc}n^{b}\tensor{h}{{}^{c}_{a}}=0,$ (76)
$S_{a}^{\rm Matter}$$\coloneqq$ $\tensor{h}{{}_{a}^{b}}n^{c}\nabla_{c}C^{\rm Matter}_{b}=0.$ (77)
We can study the well-posedness of the evolution equations with matter in a simple way. Since vacuum UG has a good initial value formulation, then we know that the evolution equations are quasilinear, diagonal, second-order hyperbolic (QDSOH), where one can use the celebrated Leray theorem Leray (1953). Thus, as long as $T_{ab}$ does not spoil the QDSOH nature of the evolution equations, and assuming that the matter equations of motion, Eq. (7), have a well-posed initial value formulation by themselves (constraints are maintained under evolution and evolution equations are QDSOH), we can claim that the theory has a well-posed initial value formulation. A sufficient set of conditions where Eq. (75) is a constraint and the evolution equations are QDSOH is when $T_{ab}$ does not have second or higher derivatives of the dynamical fields and the matter equations of motion have a well-posed initial value formulation. Of course, these are sufficient conditions; there could be cases that do not meet this hypothesis and still have a well-posed initial value formulation.
IV Conclusions
One of the most important properties of any theory is its capacity to make predictions out of initial data. Therefore, if UG is to be considered as a viable physical theory, it needs to have a well-posed initial value formulation. Moreover, to utilize this theory, we need to find all constraints required to have a consistent evolution. Fortunately, we were able to show that this theory is well-posed and to find all the constraints. For the latter, we followed the well-known Dirac method, which, produced primary constraints, $C_{a}=0$, and secondary constraints, $S_{a}=0$. Then, we used these constraints in the evolution equations to cast them in the form known from GR, where the evolution is known to have all the properties required to claim that the theory has a well-posed initial value formulation, completing the proposed analysis.
We find it remarkable that the unimodular constraint does not behave as a constraint in the sense that it is not imposed on the initial data and is preserved under evolution. This feature ought to be present in any theory with nondynamical fields like the gravitational sector of the Standard Model Extension Kostelecký (2004). Still, we show that the unimodular constraint can always be satisfied by a choice of the lapse function. This sheds light on the interpretation of the nondynamical volume and helps explain why, in UG, we can work with the traceless Einstein equations without considering the unimodular constraints.
We also found it interesting that the equivalence between GR and vacuum UG when $Q$ is constant is not a priori obvious in an initial value formulation. Interestingly, only when requiring dynamical consistency, does a spacetime constant emerge (e.g., $R=4\Lambda_{0}$ in vacuum), which can be used to show this dynamical equivalence. Still, UG can be an interesting test theory to perform numerical computation of modified geometrical gravity theories, even when $Q$ is constant. Of course, the most interesting situation is to consider matter with a variable $Q$, whose evolution is not allowed in GR.
What is more, UG has more constraints than the $4$ constraints of GR. This enhanced number of constraints is related to the reduced symmetries Henneaux and Teitelboim (1992), but a detailed counting of constraints is not straightforward Henneaux and Teitelboim (1989) and it should be done using Hamiltonian methods, which have been used, for example, to consider the problem of time in UG Unruh (1989); Smolin (2009). Of course, a full Hamiltonian analysis of the constraints could be extremely interesting. However, we expect, based on the experience with other theories with nondynamical fields Reyes and Schreck (2021); Bonder (2023), that such a study will be complicated, and it was thus left out of the present paper.
Acknowledgements.
We acknowledge valuable interactions with Néstor Ortiz, Marcelo Salgado, and Daniel Sudarsky. Also, this project benefited from financial support from CONACyT through the grant FORDECYT-PRONACES 140630 and from the UNAM-DGAPA-PAPIIT grant IN101724.